OpenStudy (anonymous):

What is the mass of 5.00 L of whole blood? densiry of blood is 1.05g/cm^3

OpenStudy (anonymous):

\[5.00L(\frac{1000ml}{1 L})(\frac{1.05 g}{1 ml})=5250g\]

OpenStudy (anonymous):

1ml is equal to 1 cm^3

OpenStudy (anonymous):

Strange i wrote out a factor label solution but it's not loading. Let me rewrite it

OpenStudy (anonymous):

5.00 L (1000 ml/ 1 L)(1.05 g/ 1 ml)=5250g